OpenStudy (anonymous):

Find the derivative of f(x) tan 3x - csc 4x.

OpenStudy (anonymous):

So you mean\[\frac{ d }{ dx }(\tan3x-4x)*\frac{ d }{ dx }(3x)*\frac{ d }{ dx }(4x)\]?

OpenStudy (anonymous):

no the derivative of tan 3x - 4x...the multiplying part is the directions

OpenStudy (anonymous):

hold on wait its tan 3x - csc 4x

OpenStudy (anonymous):

its tan 3x- csc 4x i miss typed it

OpenStudy (anonymous):

That's the answer you got or the correct answer?

OpenStudy (anonymous):

And what part of the equation is inside the tan? Is it tan(3x)-4x or tan(3x-4x)

OpenStudy (anonymous):

No not the correct answer i need to find the derivative of that and it says to multiply the derivatives of 3x and 4x into the promblem

OpenStudy (anonymous):

its tan 3x - csc 4x thats how it written

OpenStudy (anonymous):

@onegirl Please, post the right question. I confuse, too

OpenStudy (anonymous):

okay hold on

OpenStudy (anonymous):

let me edit my original question

OpenStudy (anonymous):

there do you guys see it now?

OpenStudy (anonymous):

yes. i think Twis7ed will guide you well.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Oh, ok, so\[\frac{ d }{ dx }(\tan 3x - \csc 4x)=\frac{ d }{ dx }\tan3x - \frac{ d }{ dx }\csc4x\] From that you know that \[\frac{ d }{ dx }\tan3x = 3\sec^2(3x)\] and that \[\frac{ d }{ dx }\csc4x=-4\cot(4x)\csc(4x)\]so you end up with\[\frac{ d }{ dx }(\tan3x-\csc4x)=3\sec^2(3x)+4\cot(4x)\csc(4x)\]

OpenStudy (anonymous):

^Do you understand that?

OpenStudy (anonymous):

it's the best

OpenStudy (anonymous):

yes, so i dont have to multiply the derivatives of 3x and 4x into the whole problem?

OpenStudy (anonymous):

yes i understood what you wrote/

OpenStudy (anonymous):

I'm not sure, what is the entire question?

OpenStudy (anonymous):

Okay so I found the derivative of that and my teacher told me to multiply 3x and 4x into it here is how i did it hold on

OpenStudy (anonymous):

d/dx (tan(3)x - csc 4x = tan(3) - csc 4

OpenStudy (anonymous):

i think i understnad where i went wrong now after seeing what you showed

OpenStudy (anonymous):

ALright, just try to remember that when you have a trig function you usually use the chain rule with the function and then the function inside of it and it should be easy :)

OpenStudy (anonymous):

okay thanks